OpenStudy (anonymous):

What is the arc length if Θ = 2 pi over 7 and the radius is 3 cm?

OpenStudy (anonymous):

7 pi over 12 12 pi over 7 6 pi over 7 7 pi over 6

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well isn't arc length \[l = \theta \times r\] so multiply that angle by the radius

OpenStudy (anonymous):

How is it not arc length? So, it would be L= 2pi/7 x 3?

OpenStudy (campbell_st):

well I said..."isn't arc length"

OpenStudy (campbell_st):

that's correct... you just need to simplify it

OpenStudy (anonymous):

yeah, im not really good at that :( sorry, it never caught on

OpenStudy (campbell_st):

thats ok.. so what answer do you think...?

OpenStudy (anonymous):

I guessing 6pi/7 ?

OpenStudy (campbell_st):

great guess

OpenStudy (anonymous):

lol, yaay! thank you

OpenStudy (campbell_st):

glad to help

OpenStudy (anonymous):

Question though, how do you solve that? Like changing a radian into a degree and vice versa?

OpenStudy (anonymous):

assuming the angle is in radians.... you should say so. r =3cm, diameter = 6 cm and circum = 6π arc length = circum(angle/2π) .......use this formula to get your answer...........

OpenStudy (campbell_st):

ok... so you need to know \[180^o = \pi\] so \[1^o = \frac{\pi}{180}\] and with radians \[1^r = \frac{180}{\pi}\] so as a simple explanation... an equalateral triangle has all sides equation and all angles 60 imagine a sector when the 2 radii and the arc are the same length... its like pushing on 2 vertices of an equlilateral and bending the 3rd side to make an arc... the angle goes from 60 to a value a bit less than 60... and its called a radian... so 1 radian = 57.29 degrees.... so by bending the side out... the angle has to get smaller.. hope that's not 2 confusing

OpenStudy (anonymous):

kind of, yeah.

OpenStudy (campbell_st):

well just remember 180 = pi.... and then just manipulate that

OpenStudy (anonymous):

Ok, so if I have 90 degrees and want to change it into a radian I would use 180/pi?

OpenStudy (campbell_st):

so divide both sides by 2 \[\frac{180}{2} = \frac{\pi}{2} ...or.... 90^o = \frac{\pi}{2}\]

OpenStudy (anonymous):

45?

OpenStudy (campbell_st):

ok.... so 45 is half of 90.. so divide the previous by 2 \[\frac{90}{2} = \frac{\frac{\pi}{2}}{2}.... or...... 45 = \frac{\pi}{4}\]

OpenStudy (anonymous):

11.25?

OpenStudy (anonymous):

wait, is pi/4 the answer?

OpenStudy (campbell_st):

thats it... where ever possible leave the angle in terms of pi

OpenStudy (anonymous):

oh ok! Thanks for explaining, I'm a little rusty but I think I get it

OpenStudy (anonymous):

Can you help with another question? It's about amplitude and periods etc

OpenStudy (campbell_st):

ok... just post it... as a new question... and there are plenty of people who can help

OpenStudy (anonymous):

ok thanks.

OpenStudy (anonymous):

Thanku soo much!! :D

OpenStudy (pink.1d):

@malice ;)

OpenStudy (pink.1d):

@malice this is how i found u hehe

OpenStudy (anonymous):

LOL cool haha :D

OpenStudy (anonymous):

I can't believe that I have already finished this course... *sniff* ^-^

OpenStudy (anonymous):

Life really does go on.............

OpenStudy (pink.1d):

@malice im still finishing it